OpenStudy (alfie):

Sequence of functions and uniform convergence.. I am having some troubles with this topic, so here it is... \[f_{n}(x)\] = \[\large \log(1+\frac{1}{n(x-1)})\] Now.. Limit for n->infinity is log1 so 0. So it converges pointwise to 0. Now I have to find out if it uniformly converges, so here it is: \[\large \sup|f_{n}(x) - f(x)| = \sup|f_{n}(x)|\] Should go to 0. Now I start to get confused. What I have to do exactly, take the first derivative respect to x and = 0 ? Help!

OpenStudy (anonymous):

x in R?

OpenStudy (alfie):

sorry, x belonging to ]1,+infinity[

OpenStudy (anonymous):

Ok, you must be prove the following inequality : \[\ln(1+u) \le u, \forall u \in (0,1)\]

OpenStudy (alfie):

I am not sure my calculus teacher actually used this method, you got any reference?

OpenStudy (anonymous):

because we want to have a result : \[\sup_{x>1} |f_n (x)| \to 0\], and we see that \[\frac{1}{n} \to 0\] so we need take \[\frac{1}{n(x-1)}\] out of the log. By using this inequality, sorry, we can got it, but may be we are entangled because of (x-1)

OpenStudy (anonymous):

if x>2, then got it :D

OpenStudy (anonymous):

what interval are you on?

OpenStudy (anonymous):

oh nvm i see it, then you have a problem. it is clear what you have to do to prove uninform convergence or no? you have to show that this limit is independent of x, i.e. it depends only on n it would be good to negate the definition of uniform convergence and see exactly what it says. then i believe you should be able to prove that this does not converge uniformly

OpenStudy (anonymous):

do you know how to negate the definition?

OpenStudy (anonymous):

in fact the proof that it is not uniform is pretty straightforward once you write what it is you have to show

OpenStudy (anonymous):

@satellite73 yes it is. This sequence does not converge uniformly. Concemtrate your thinking close to 1.

OpenStudy (anonymous):

a hint would be to find \(x\) such that given any \(n\) you can make it larger than \(\log(2)\)

OpenStudy (anonymous):

think @alfie went away though the important part of the problem is not the proof, but rather understanding how to negate a definition.

OpenStudy (anonymous):

If it is uniformly convergent, then there exits N, such that n>N one has \[ \ln \left( 1 + \frac 1 {n(x-1)} \right) \le \frac 1 2 \] for every x Take \[ x_n = 1 +\frac 1{n^2} \] in the above inequality and find a contradiction.

OpenStudy (anonymous):

If we do that we get for every n >N \[ \ln ( 1 + n) < 1/2 \] Which cannot be.

OpenStudy (alfie):

okay our calculus professor didn't really explain it this way but sounds good to me. Thank you guys, if I could have given more best answers I'd have done that. Thanks ;)

OpenStudy (mendicant_bias):

I'm failing to understand why it is pointwise convergent; I've been struggling with this for a little in general, but I don't see explicitly how M is dependent on Epsilon, could somebody write it out or something?